OpenStudy (anonymous):

Evaluate lim x^2+x ---- ------- x->-1 x^2-1

OpenStudy (anonymous):

factor and cancel

OpenStudy (anonymous):

\[\frac{x^2+x}{x^2-1}=\frac{x(x+1)}{(x+1)(x-1)}=\frac{x}{x-1}\]

OpenStudy (anonymous):

then replace x by -1 to get \[\frac{1}{2}\]